OpenStudy (sloppycanada):

Find f(g(x)) if f(x) = 2x + 1 and g(x) = x2 - 1.

OpenStudy (jhannybean):

Wherever you see an x in your function for f(x) replace it with the function of g(x). \[f(g(x)) = 2(x^2-2)+1\]

OpenStudy (anonymous):

did u understood?

OpenStudy (sloppycanada):

\[2x^2-3\] @Jhannybean

OpenStudy (sloppycanada):

but that's not one of my choices.

OpenStudy (jhannybean):

What are your answer choices?

OpenStudy (anonymous):

ok \[f(g(x))= 2(x ^{2}-1)+1\]

OpenStudy (anonymous):

\[f(g(x))=2x ^{2}-2+1\]

OpenStudy (anonymous):

2(x^2-1)+1 = 2x^2 -2 + 1 = 2x^2 - 1

OpenStudy (jhannybean):

Oh, I made a typo there.

OpenStudy (jhannybean):

The \(x^2-2\) should have been \(x^2-1\)

OpenStudy (anonymous):

\[f(g(x))= 2x ^{2}-1\]

OpenStudy (anonymous):

hope this helps :)

OpenStudy (jhannybean):

No harm in correcting me :)

OpenStudy (anonymous):

f(g(x)) = 2(g(x)) +1 = 2(x^2-1)+1

OpenStudy (sloppycanada):

So this one - Find g(f(x)) if f(x) = 2x - 3 and g(x) = x - 1.

OpenStudy (sloppycanada):

Should be - \[2x^2 + x -3\]

OpenStudy (anonymous):

g)x)-1 = (2x-3)-1 = 2x-4

OpenStudy (sloppycanada):

Find f(g(x)) if f(x) = 2x + 1 and g(x) = x2 - 1. My answer options are - 2x -1 \[2x^2-1\] \[2x^2 + 1\] \[2x^3-1\]

OpenStudy (sloppycanada):

So if it's f(g(x)), shouldn't it be something along the lines of - (x^2 - 1)(2x+1)?

OpenStudy (jhannybean):

We've done this one.

OpenStudy (jhannybean):

Just scroll up?

OpenStudy (sloppycanada):

It's just 2x -1 ?

OpenStudy (jhannybean):

You're not multiplying anything actually, you're just substituting the function of g(x).

OpenStudy (sloppycanada):

Wait what? I am very confused.

OpenStudy (jhannybean):

\[f(g(x))= f(x^2-1) = 2(x^2-1)+1 = 2x^2-2+1 =2x^2-1\]

OpenStudy (jhannybean):

Do you see it now, @sloppycanada ?

OpenStudy (sloppycanada):

Kind of..

OpenStudy (jhannybean):

which part is confusing you?

OpenStudy (sloppycanada):

Where the "x" in 2x-1 went.

OpenStudy (jhannybean):

It was replaced by g(x)

OpenStudy (sloppycanada):

Okay, so in this - Find g(f(x)) if f(x) = 2x - 3 and g(x) = x - 1. It'd just be 1(2x -3) - 1?

OpenStudy (jhannybean):

Not quite, \[g(f(x)) = g(2x-3)= (2x-3)-1 = 2x-4\]

OpenStudy (sloppycanada):

Okay so again - Find (f – g)(x) if f(x) = 2x – 6 and g(x) = 3x + 8. (2x -6) - (3x+ 8) = -x +2?

OpenStudy (jhannybean):

Somewhat, but a small math error. \[(f-g)(x) = f(x) - g(x) = (2x-6)-(3x+8) = -x-(8+6) = -x-14\]

OpenStudy (jhannybean):

You've got the idea though :D

OpenStudy (sloppycanada):

Find (f + g)(x) if f(x) = 1/2x + 4 and g(x) = -x - 3. (1/2x +4) + (-x-3) = -1/2 + 1

OpenStudy (sloppycanada):

-1/2x* +1?

OpenStudy (jhannybean):

Yep

OpenStudy (sloppycanada):

Find the zeros of -4x + 6